yoooweslime:

1. What is the GCF of: 9a^4 b^4 - 27a^3b^3 2. Factor using the difference of two square methods: x^2 - 36 3. Factor the trinomial: x^2 + 7x + 10

yoooweslime:

@kjsaif

yoooweslime:

@hhanan

yoooweslime:

@0mega

yoooweslime:

i really need help i have a test in an hour

hhanan:

okay so for the first one you have to Find the Greatest Common Factor then Factor out the GCF then simplify it So you get :\[9a^3b^3(ab-3)\] For the 2 one you have to 1 rewrite the form \[x^2-6^2\] then you get (x+6)(x-6) For the 3 one the number that adds to 10 and 7 are 2 and 5 because 5+5 is 10 and 5+2 is 7 So just rewrite the expression and you get \[(x+2)(x+5)\] I have to go so if you have any other question I might be back later

yoooweslime:

@hhanan wrote:
okay so for the first one you have to Find the Greatest Common Factor then Factor out the GCF then simplify it So you get :\[9a^3b^3(ab-3)\] For the 2 one you have to 1 rewrite the form \[x^2-6^2\] then you get (x+6)(x-6) For the 3 one the number that adds to 10 and 7 are 2 and 5 because 5+5 is 10 and 5+2 is 7 So just rewrite the expression and you get \[(x+2)(x+5)\] I have to go so if you have any other question I might be back later
Thank yu

surjithayer:

for 3 break 7 in a+b=7 \[a \times b=10\] 7=5+2 \[5 \times 2=10\] \[x^2+(5+2)x+10\] \[=x^2+5x+2x+10\] \[=x(x+5)+2(x+5)\] \[=(x+5)(x+2)\]